OpenStudy (anonymous):

solve: 36/27=2x/3 A. 4 B. 2 C. 3 D. 6

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

so you solve this kind of thing by getting x by itself. what can you do to get x by itself?

jimthompson5910 (jim_thompson5910):

36/27=2x/3 36*3=27*2x ... cross multiply ... ... ... x = ???

OpenStudy (anonymous):

108=54

OpenStudy (espex):

Not quite.

OpenStudy (anonymous):

\[\frac{36 }{ 27} = \frac{2x }{ 3 }\\\\ \frac{ 36 }{9 } = 2x \\\\\ 4 =2x\\ x=2\]

OpenStudy (anonymous):

wait im confused now what do we do after ?

OpenStudy (espex):

Well you needed to keep your 108=54 but you forgot the 'x' so it would be 108=54x. Then you divide the 54 to the other side so you can solve for x.